The Magellanic Clouds Newsletter

An electronic exchange of Magellanic Clouds information

Edited by: Eva K. Grebel, You-Hua Chu, and Dominik J. Bomans

                                 mcnews@astro.uiuc.edu
                      http://www.astro.uiuc.edu/mcnews/MCNews.html
                         http://www.astro.uni-bonn.de/~mcnews/

No. 20 May 3, 1998

Table of Contents

Abstracts of 10 refereed papers 1
Abstract of a non-refereed paper 7

Abstracts of Refereed Papers

An Hi Aperture Synthesis Mosaic of the Large Magellanic Cloud

Sungeun Kim^{1}, L. Staveley-Smith^{2}, M.A. Dopita^{1}, K.C. Freeman^{1}, R.J. Sault^{2}, M.J. Kesteven^{2}, and D. McConnell^{2}

1 Mount Stromlo and Siding Spring Observatories, Weston Creek PO. ACT 2611, Australia 2 Australia Telescope National Facility, Epping 76, NSW 2121, Australia

We present the results of an Hi aperture synthesis mosaic of the Large Magellanic Cloud (LMC), made by combining data from 1344 separate pointing centers using the Australia Telescope Compact Array^{1} (ATCA). The resolution of the mosaiced images is 1.′0 (15 pc, using a distance to the LMC of 50 kpc). This mosaic, with a spatial resolution 15 times higher than that which had been previously obtained, emphasizes the turbulent and fractal structure of the ISM on the small scale, resulting from the dynamical feedback of the star formation processes with the ISM. The structure of the neutral atomic ISM in the LMC is dominated by Hi filaments combined with numerous shells and holes. On the large scale, Hi disk appears to be remarkably symmetric, and to have a well organized and orderly, if somewhat complex, rotational field. The bulk of the Hi resides in a disk 7.3 kpc in diameter. The mass of the disk component of the LMC is 2.5 × 10^{9} M_{⊙}, and the upper limit to all mass within a radius of 4 kpc is ∼ 3.5 × 10^{9} M_{⊙}.

1The Australia Telescope is funded by the Commonwealth of Australia for operation as a National Facility managed by CSIRO.

Accepted by: The Astrophysical Journal (Aug 20) For preprints, contact sek@mso.anu.edu.au

Also available from the URL http://www.atnf.csiro.au/research/lmc h1 (after May 8)

Interstellar Abundances in the Magellanic Clouds. II. The Line of Sight to SN 1987A in the LMC

D.E. Welty^{1}, P.C. Frisch^{1}, G. Sonneborn^{2}, & D.G. York^{1}

1 University of Chicago, Astronomy & Astrophysics Center, 5640 S. Ellis Ave., Chicago, IL 60637, USA 2 Goddard Space Flight Center, Lab. for Astronomy & Solar Physics, Code 681, Greenbelt, MD 20771, USA

We have analyzed both high-resolution optical absorption-line spectra and UV spectra obtained with IUE of the LMC SN 1987A, in order to determine abundances and physical conditions in the various neutral interstellar clouds along the line of sight to the SN. We have used a flat-fielding procedure to enhance the S/N ratios and the reliability of weak features in the UV spectra, and have modeled the UV line profiles using the component structure derived from the higher resolution, high S/N ratio optical spectra of Caii and Nai. Fits to the Caii, Cai, and Nai absorption-line profiles reveal (at least) 44 components, at velocities −24 km s^{−1} ≤ v_{⊙} ≤ 295 km s^{−1}, which can be associated with the ten component groups discernible in the lower resolution UV spectra. From the UV spectra, we determined component group column densities for Ci, Mgi, Mgii, Alii, Siii, Pii, Cli, Tiii, Crii, Mnii, Feii, Niii, and Znii — with 1σ uncertainties less than 0.1 dex in many cases. These are the most extensive and accurate abundances yet measured for the neutral ISM in the LMC.

The component velocities, the patterns of relative elemental abundances [X/Zn] and [X/Fe], and various diagnostic ratios have been used to estimate the locations and physical characteristics [N(H), T, n_{e}] of these component groups. (Systematic differences among the diagnostic ratios make the derived physical properties somewhat uncertain, however.) The components at low velocities (5 km s^{−1} ≤ v ≤ 23 km s^{−1}) have relative abundances and values for the diagnostic ratios very similar to those found for warm, diffuse Galactic disk clouds, and likely are due to a mixture of warm and cool gas in the Galactic disk. The components at velocities 56 km s^{−1} ≤ v ≤ 91 km s^{−1} are due to a mixture of warm and cool gas, apparently with negligible depletions, in the Galactic halo. The two intermediate velocity component groups (109 km s^{−1} ≤ v ≤ 138 km s^{−1} and 155 km s^{−1} ≤ v ≤ 178 km s^{−1}) both have relative abundances similar to those found for Galactic halo clouds. These warm (T ≥ 4500 K), partially ionized clouds are probably located in the Galactic halo and in the LMC, respectively. The components at velocities 193 km s^{−1} ≤ v ≤ 225 km s^{−1} also have relative abundances similar to those in the halo clouds, but are likely due to relatively warm gas in the LMC, perhaps very close to the SN. The component groups at 239 km s^{−1} ≤ v ≤ 256 km s^{−1} and 265 km s^{−1} ≤ v ≤ 270 km s^{−1} are probably located on opposite sides of the main LMC component group (at velocities 275 km s^{−1} ≤ v ≤ 296 km s^{−1}) (using absorption-line data for several other adjacent lines of sight and the structure inferred from SN light echo observations). The relative abundances and diagnostic ratios for those three LMC groups are similar to those found for warm, low density Galactic disk clouds. Higher Nai/Caii ratios, the presence of CH, and the Ci fine structure level populations suggest, however, that the main LMC group contains both cool and warm gas — consistent with the results from several recent 21 cm emission-absorption studies of the LMC.

Since the abundance patterns determined for stars and gaseous nebulae in both the SMC and the LMC appear to be similar to the solar pattern (for the elements whose interstellar abundances we have considered), the similarities in relative gas phase interstellar abundances in our Galaxy and in the MC suggest that the dust depletions follow similar patterns as well — despite differences in metallicity and dust-to-gas ratio among the three galaxies. These local relative abundance/depletion patterns can thus be used to infer total (gas + dust) abundances for QSO absorption-line systems at various redshifts.

Submitted to: The Astrophysical Journal

For preprints, contact welty@oddjob.uchicago.edu

Light variations of massive stars (α Cyg variables) XV. The LMC supergiants R 99 (LBV), R 103, R 123 (LBV) and R 128

A.M. van Genderen^{1}, C. Sterken^{2}, M. de Groot^{3} and R.A. Reijns^{1}

1 Leiden Observatory, Postbus 9513, NL-2300RA Leiden, The Netherlands 2 University of Brussels (VUB), Pleinlaan 2, B-1050 Brussels, Belgium 3 Armagh Observatory, College Hill, Armagh BT61 9DG, Northern Ireland, UK

V BLUW photometry (Walraven system) of the four variable LMC supergiants R 99, R 103, R 123 and R 128 is analysed, searched for periods and discussed. Based on former and present photometry we conclude that two of the three emission-line objects are undoubtedly active LBVs (R 99 and R 123), although not so spectacular. R 123, like AG Car near minimum brightness, shows a low amplitude S Dor activity with superimposed α Cyg-type variations.

Accepted by: Astronomy & Astrophysics

For preprints, contact genderen@strw.leidenuniv.nl

ROSAT archival observations of 1SAX J0054.9–7226 = 2E 0053.2–7242, a newly discovered X–ray pulsar in the SMC

G.L. Israel^{1}, S. Campana^{2}, G. Cusumano^{3}, F. Frontera^{4}, M. Orlandini^{4}, A. Santangelo^{3} and L. Stella^{1}

1 Osservatorio Astronomico di Roma, Via Frascati 33, I–00040 Monteporzio Catone (Roma), Italy 2 Osservatorio Astronomico di Brera, Via E. Bianchi 46, I–23807 Merate (Lecco), Italy 3 Istituto di Fisica Cosmica e Applicazioni all’Informatica, C.N.R., Via La Malfa 153, I–90146 Palermo, Italy 4 Istituto Tecnologie e Studio Radiazioni Extraterrestri, C.N.R., Via Gobetti 101, I–40129 Bologna, Italy

We analysed 13 archival ROSAT PSPC and HRI observations which included the position of a newly discovered 59 s X–ray pulsar in the Small Magellanic Cloud, 1SAX J0054.9–7226 = 2E 0053.2– 7242. The source was detected three times between 1991 and 1996 at a luminosity level ranging from ∼8×10^{34} - 4× 10^{35} erg s^{−1} (0.1–2.4 keV). Highly significant pulsations at 59.072 s were detected during the 1991 October 8–9 observation. The ROSAT period, together with those measured by RXTE and BeppoSAX yields a period derivative of Ṗ= – 0.016 s yr^{−1}. A much more accurate source position (10′′ uncertainty) was obtained through the ROSAT HRI detection on 1996 April restricting to three m_{V} > 15.5 stars the likely counterpart of 1SAX J0054.9–7226 = 2E 0053.2–7242. Future optical follow–up observations of these candidates should determine the counterpart of 2E 0053.2–7242 and its probable Be star X–ray transient nature. The optical and/or infrared activity brightening of the counterpart will allow further X–ray triggers and studies.

Accepted by: Astronomy & Astrophysics Letters

For preprints, contact gianluca@coma.mporzio.astro.it
Also available from the URL http://xxx.lanl.gov/ps/astro-ph/9804186

The optical counterparts to Be/X-ray binaries in the Magellanic Clouds

J.B. Stevens^{1}, M.J. Coe^{1} and D.A.H. Buckley^{2}

1 Dept. of Physics and Astronomy, University of Southampton, Highfield, Southampton, SO17 1BJ, UK 2 South African Astronomical Observatory, P.O. Box 9, Observatory, 7935, South Africa

The fields of 8 X-ray sources in the Magellanic Clouds believed to be Be/X-ray binaries have been searched for possible Be star counterparts. BV R and Hα CCD imaging was employed to identify early type emission stars through colour indices and Hα fluxes. Follow up Hα spectroscopy of 5 sources confirms the presence of Hα emission in each case. Based on the positional coincidence of emission line objects with the X-ray sources, we identify Be star counterparts to the ROSAT sources RX J0032.9-7348, RX J0049.1-7250, RX J0054.9-7226 and RX J0101.0-7206, and to the recently discovered ASCA source AX J0051-722. We confirm the Be star nature of the counterparts to the EXOSAT source EXO0531.1-6609, and the HEAO1 source H0544-665. In the field of the ROSAT source RX J0051.8-7231 we find three possible counterparts, each showing evidence for Hα emission.

Submitted to: Monthly Notices of the Royal Astronomical Society

For preprints, contact jbs@astro.soton.ac.uk

The possible effects of an unusual resonance in very long period Cepheids

E. Antonello^{1}

1 Osservatorio Astronomico di Brera, Via E. Bianchi 46, 23807 Merate, Italy

The shape of the radial velocity and light curves of 24 long-period (30 ≤ P ≤ 134 d) Cepheids in the Magellanic Clouds shows a progression with the period. The sequences of the radial velocity and light curves are based only on a small sample of stars; however, evident changes of the shape can be seen in Cepheids with period between 90 and 134 d. The Fourier parameter–period diagrams for the radial velocity curves show trends which remind in part those of Cepheids with period near 10 d. The plausible interpretation is a resonance, probably P_{0}/P_{1} = 2 between the fundamental and the first overtone mode. The possible importance of this phenomenon for the study of stellar structure and evolution in relatively far galaxies is emphasized.

Accepted by: Astronomy & Astrophysics Letters

For preprints, contact elio@merate.mi.astro.it
Also available from the URL http://www.merate.mi.astro.it/~elio/elio.html

Cepheid models based on self-consistent stellar evolution and pulsation calculations: the right answer?

I. Baraffe^{1}, Y. Alibert^{1}, D. Méra^{1}, G. Chabrier^{1}, and J.P. Beaulieu^{2}

1 C.R.A.L, Ecole Normale Supérieure, 69364 Lyon Cedex 07, France 2 Kapteyn Institute, PO box 800, 9700 AV Groningen, The Netherlands

We have computed stellar evolutionary models for stars in a mass range characteristic of Cepheid variables (3 < m/M_{⊙} < 12) for different metallicities representative of the Galaxy and the Magellanic Clouds populations. The stellar evolution calculations are coupled to a linear non adiabatic stability analysis to get self-consistent mass-period-luminosity relations. The period - luminosity relation as a function of metallicity is analysed and compared to the recent EROS observations in the Magellanic Clouds. The models reproduce the observed width of the instability strips for the SMC and LMC. We determine a statistical P-L relationship, taking into account the evolutionary timescales and a mass distribution given by a Salpeter mass function. Excellent agreement is found with the SMC PL relationship determined by Sasselov et al. (1997). The models reproduce the change of slope in the P-L relationship near P ∼ 2.5 days discovered recently by the EROS collaboration (Bauer 1997; Bauer et al. 1998) and thus explain this feature in term of stellar evolution. Some discrepancy, however, remains for the LMC Cepheids.

The models are also in good agreement with Beat Cepheids observed by the MACHO and EROS collaborations. We show that most of the 1H/2H Beat Cepheids have not yet ignited central helium burning; they are just evolving off the Main Sequence toward the red giant branch.

Accepted by: The Astrophysical Journal Letters

For preprints, contact ibaraffe@ens-lyon.fr
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9804061

Mass Segregation in two young clusters in the Large Magellanic Cloud: SL 666 and NGC 2098

M. Kontizas^{1}, D. Hatzidimitriou^{2}, I. Bellas-Velidis^{3}, D. Gouliermis^{1,3}, E. Kontizas^{3}, and R.D. Cannon^{4}

1 Section of Astrophysics Astronomy & Mechanics, Department of Physics, University of Athens, GR-157 83 Athens, Greece 2 Physics Department, University of Crete, P.O. Box 2208, Heraklion, 710 03, Crete, Greece 3 Astronomical Institute, National Observatory of Athens, P.O. Box 20048, GR-118 10, Greece 4 Anglo-Australian Observatory, Epping Laboratory, P.O. Box 296, Epping, NSW2121, Australia

The age and dynamics of the young LMC clusters SL 666 and NGC 2098 were investigated using CCD observations obtained with the 3.9m Anglo-Australian Telescope. SL 666 was found to have an age of 1 − 1.25 × 10^{8} yr, while NGC 2098 is somewhat younger, with an age of 6.3 − 7.9 × 10^{7} yr. In the case of SL 666 the good quality of the acquired data allowed the construction of the main sequence luminosity function as a function of radius, which revealed mass segregation, also confirmed from star counts. Since the age of SL 666 is significantly lower than the mean relaxation time of at least 5 × 10^{8} yr calculated for this cluster and since the mass range of the cluster members does not permit us to assume faster energy equipartition, the observed mass segregation seems to be due to the way stars form in the cluster. Indications on a star formation scenario are examined. Disruption time is also discussed and there is strong evidence that the cluster is bound. NGC 2098 shows strong evidence of mass segregation, but further observations are needed in this case.

Accepted by: Astronomy & Astrophysics For preprints, contact mkonti@astro.noa.gr

Three Populous Clusters Discovered to be in the LMC Age Gap

Ata Sarajedini

1 San Francisco State University, 1600 Holloway Avenue, San Francisco, CA 94132, USA

In the process of searching the Hubble Space Telescope archive, we have serendipitously discovered three populous Large Magellanic Cloud (LMC) clusters with ages that place them in the LMC ‘age gap’. These clusters – NGC 2155, SL 663, and NGC 2121 – turn out to have [Fe/H] ≈ −1.0 and ages of ≈ 4 Gyr. This puts them in the age gap between the intermediate-age LMC clusters, the oldest of which are ≈ 2.5 Gyr old, and ESO121-SC03, which has an age of ≈ 9 Gyr. The addition of these three clusters to the LMC age - metallicity relation has reduced the discrepancy between the age distribution of the LMC clusters and the field stars. Furthermore, it indicates that searches to find more clusters older than ≈ 2.5 Gyr in the LMC are crucial to a better understanding of its global star formation history.

Accepted by: The Astronomical Journal For preprints, contact ata@ucolick.org

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9804340

Age, Metallicity, and the Distance to the Magellanic Clouds From Red Clump Stars

Andrew A. Cole^{1}

1 University of Wisconsin-Madison, 475 N. Charter St., Madison, WI, 53706-1582, USA

We show that the luminosity dependence of the red clump stars on age and metallicity can cause a difference of up to < ∼0.6 mag in the mean absolute I magnitude of the red clump between different stellar populations. We show that this effect may resolve the apparent ≈0.4 mag discrepancy between red clump-derived distance moduli to the Magellanic Clouds and those from, e.g., Cepheid variables. Taking into account the population effects on red clump luminosity, we determine a distance modulus to the LMC of 18.36 ± 0.17 mag, and to the SMC of 18.82 ± 0.20 mag. Our alternate red clump LMC distance is consistent with the value (m-M)_{LMC} = 18.50 ± 0.10 adopted by the HST Cepheid Key Project. We briefly examine model predictions of red clump luminosity, and find that variations in helium abundance and core mass could bring the Clouds closer by some 0.10–0.15 mag, but not by the ≈0.4 mag that would result from setting the mean absolute I-magnitude of the Cloud red clumps equal to the that of the Solar neighborhood red clump.

Accepted by: The Astrophysical Journal Letters For preprints, contact cole@astro.wisc.edu

Also available from the URL www.astro.wisc.edu/~cole

Abstracts of Non-Refereed Papers

Ultraviolet Observations of Stars in The Magellanic Clouds: A Historical Bridge from the Sun to R 136 and Beyond.

J. Wm. Parker^{1}

1 Southwest Research Institute, Suite 426, 1050 Walnut Street, Boulder, CO, 80302, USA

The Magellanic Clouds have provided us with unique insights into the lives of massive stars and have served as test beds for star formation and evolutionary scenarios and for constraining stellar parameters. From an observational point of view, the LMC and SMC have a number of advantages over the Galaxy, with an almost ideal combination of comparatively small (but variable) line-of-sight reddening, different metallicities and reddening laws, rich populations of massive stars, a relatively well-known distance for each galaxy, and such a distance that allows us to make both global and smallscale observations of the stellar populations and their environments. In most cases, these properties make the Clouds ideal targets for UV observations.

I review the last few decades of UV observations of the Magellanic Clouds, and how these data have been instrumental in advancing our knowledge of massive stars. In particular, I present observations made by the Ultraviolet Imaging Telescope (UIT ), which obtained photometric data over a wide field of view (images are 37 arcmin in diameter). These data are excellent at highlighting hot stars while suppressing the cooler, lower mass stars that often cause confusion in studies of crowded fields. Using the UIT data, I derive initial mass functions and other parameters for field stars as well as for stars in a collection of clusters. I discuss other uses of these UV data and their utility for identifying populations of massive stars in the Clouds for future studies.

To appear in: Hot Stars in Open Clusters of the Galaxy and the Magellanic Clouds, RMxAA Conference Series, (ed. V. Niemela & N. Morrell).

For preprints, contact joel@boulder.swri.edu
Also available from the URL http://www.boulder.swri.edu/~joel/papers.html
or by anonymous ftp at ftp.boulder.swri.edu/pub/joel/uv clouds.ps.gz